Find the solution to the initial value problem 1 [E]-: :] [:] [:O] - [8]
Answer #1

Doubt in this then comment below.. i will help you..

.

please thumbs up for this solution.. thanks..

.

xs AX | A-Il to -> 3 6-1 2 -7 7+120 d 4,3 for dry (A-4I) Xro - 2 13 3 2) y = { 2 -213 fo =3 (A-3I) - ( ) x = (%) Ona () 3t yxlola () ~23 G-C2 = 9 Gt (2=5 (6 27 ( 2-22 xalta 4t - 18e 3+ +22e lat 3t X₂ ( tla -22e 2re